3x-2y=-31
5x+-2y=-49
Solve linear systems by adding or subtracting
From Holt McDougal text: Florida Larson Algebra 1
Chapter/Section 7.3 extra practice page 908 problem 15

Answer by stanbon(75887)

3x-2y=-31
5x+-2y=-49
-------------
Subtract and solve for "x":
2x = -18
x = -9
-----
Solve for "y":
3x-2y = 31
3(-9) - 2y = -31
-27 -2y = -31
2y = 4
y = 2
------
Cheers,
Stan H.
==========